OpenStudy (anonymous):

A science fair poster is a rectangle 3 feet long and 2 feet wide. What is the area of the poster in square inches?

OpenStudy (anonymous):

3 feet equal to 12*3=36 inches 2 feet equal to 12*2=24 inches 36*24= 864 inches

OpenStudy (anonymous):

the exponent in the unit of the answer is incorrect

OpenStudy (anonymous):

would it be in*2 or in*3

OpenStudy (anonymous):

in*2 is the correct answer but great math @plohrb

OpenStudy (anonymous):

yeah forgot the square inches my bad

OpenStudy (anonymous):

no your totally fine lets try another one my friend

OpenStudy (anonymous):

A rectangular floor is 5 yards long and 4 yards wide. What is the area of the floor in square feet?

OpenStudy (anonymous):

@plohrb

OpenStudy (anonymous):

keep in mind Ft, Ft^2 or Ft^3

OpenStudy (anonymous):

is this homework for you?

OpenStudy (anonymous):

No its a program by the name of ALEPT i am running a research out of the university of central Florida

OpenStudy (anonymous):

Your name will be completely anonymous

OpenStudy (anonymous):

We are trying to figure out a more secure way of learning other then FLVS

OpenStudy (anonymous):

honestly its not that hard to google either lol

OpenStudy (anonymous):

We are getting away from the topic,

OpenStudy (anonymous):

but i appreciate your service